Following macroeconomic data is given for a hypothetical economy. All Values are given in


Rs. Millions.


a) Compute Marginal Propensity to Save



b) Derive the saving function of this economy



c) Calculate the equilibrium level of national income.



d) If government purchases rise to Rs. 250 million, calculate the expenditure


multiplier and new equilibrium national income.



Consumption (C) 300 + 0.8Yd


Investment (I) 200


Taxes (Te) 100


Government transfer to household (Re) 80


Government Purchases (G) 150

1
Expert's answer
2021-11-01T17:55:43-0400

a) Marginal Propensity to Save (MPS) = "\frac{change\ in\ savings}{change\ in\ disposable}" = "MPS = \frac{ΔS}{ ΔY}"

b)The saving function "S = -a + (1 \u2013 b) (Y \u2013 T)"

c)Equilibrium level of national income is obtained when the aggregate supply (AS)= aggregate demand (AD).

d) Multiplier = "\frac{1}{1 - MPC}"


Kindly, provide more data and information to this question